Theorem ork- The Central Limit A leading magazine(like Barrons) reported at one time that the average number of weeks an individual is unemployed is 34 weeks. Assume that for the population of all unemployed individuals the population mean length of unemployment is 34 weeks and that the population standard deviation is 3.5 weeks. Suppose you would like to select a random sample of 55 unemployed individuals for a follow-up study Find the probability that a single randomly selected value is greater than 34.3. Px>34.3)(Enter your answers as numbers accurate to 4 decimal places.) Find the probability that a sample of size n # 55 is randomly selected with a mean greater than 34.3. PM 343)Enter your answers as numbers accurate to 4 decimal places.) Points possible:1 Unlimited attempts.
naniat The Central Limit Theorenm A leading magazine(like Barrons) reported at one time that the average number of weeks an individual is unemployed is 20.4 weeks. Assume that for the population of all unemployed individuals the population mean leagh of unemployment is 20.4 wecks and that the population standard deviation is 10 weeks. Suppose you would like to select a random sample of 11 unemployed individuals for a follow-up study Find the probability that a single randomly selected value is between 11.7 and 21.6. P(l1.7<X<21.6)- Find the probablity thata sample of size n P1.7<M<21.6)- - 11 is randomly selected with a mean between 11.7 and 21.6. Enter your answers as numbers accurate to 4 decimal places
Answer #1

1)

Here, X ~ N(34.3.5 )

P(X > 34.3)-P(z> 1-0.5359 0.4641 34.3-34 3.5 )-P(Z > 0.09) 1-P(Z < 0.09)

Also, 3.52

34.3-34 P(M > 34.3) = P(M > ) = P(Z > 0.64) = 1-P(Z < 0.64) 3.52 1-0.7389-0.2611

2)

Here, X ~ N(20.4. 102)

11.7 - 20.4 10 21.6- 20.4 10 P(11.7< X< 21.6)-PP(-0.87 < Z < 0.12) = P(Z < 0.12)-P(Z <-0.87)

=P(Z<0.12)-P(Z>0.87)=P(Z<0.12)-[1-P(Z<0.87)]

0.5478-11-0.8078| 0.3556

Also, Msim N(mu,rac{sigma^{2}}{n})sim N(20.4,rac{10^{2}}{11})

P(11.7<M<21.6)=P(rac{11.7-20.4}{sqrt{rac{10^{2}}{11}}}<Z<rac{21.6-20.4}{sqrt{rac{10^{2}}{11}}})

=P(-2.89<Z<0.40)=P(Z<0.40)-P(Z<-2.89)

=P(Z<0.40)-P(Z>2.89)=P(Z<0.40)-[1-P(Z<2.89)]

=0.6554-[1-0.9981]=0.6535